### v1

@@ 6 Concluding Remarks
This paper studies the intergenerational transmission of verifiable information in an overlapping generations model that incorporates both physical communication frictions as well as agents’ strategic incentives to conceal unfavorable information. We show that lowering the communication frictions will make the agents more selective in disclosing information. This is because disclosing an additional good signal has two effects on the next agent’s belief: a direct effect which shows the existence of another good signal as well as an indirect effect which shows the existence of another subsequence of bad signals. As communication frictions vanish, the expected length of each bad subsequence increases. Hence, the indirect effect will become stronger, so that fewer signals can have high enough likelihood ratios to overcome the more pronounced indirect effect. The indirect effect is absent in Dye (1985)’s model where increasing the probability that the sender has information makes his disclosure less selective.
A natural question is whether our conclusions extend to a one-shot disclosure game between a sender and a receiver, where the number of signals the sender has follows a geometric distribution with parameter $\delta$. For equilibria in constant threshold strategies considered in Theorem 2, our conclusion that agents become more selective in disclosing information as $\delta$ increases will extend. However, it is unclear whether Theorem 1 extends to this one-shot game, namely, when the equilibrium disclosure threshold could vary across histories and private signals or when the restriction on the equilibrium strategy is replaced with one on the belief map.
Intuitively, our overlapping generations model has a gradualism property stated in Section 4 that makes the analysis tractable. It leads to Lemma 1 and Lemma 2, which makes it tractable to compute the change in the receiver’s belief once the sender discloses one fewer signal $s_{j}$, the signal with the lowest likelihood ratio among those being disclosed.
In contrast, more possibilities could arise in the one-shot disclosure game, which makes the change in the receiver’s belief harder to compute. For example, it could be the case that agents never disclose $s_{2}$ when there is only one such signal available but will disclose all $s_{2}$ when at least two such signals are available. This can be sustained by the off-path belief that upon observing only one $s_{2}$, the receiver believes that the sender has hidden many signals with low likelihood ratios but will not entertain this adverse belief upon observing no $s_{2}$ or two or more $s_{2}$.
Such a possibility is ruled out in our overlapping generations model since without any agent disclosing the first realization of $s_{2}$ at histories that do not contain any $s_{2}$, no agent will have two $s_{2}$ available to disclose. Understanding the connections between our overlapping generations model and the one-shot disclosure game is an important question left for future research.
Appendix A Proof of Proposition 1
@@
If $L_{i}>1$, then the value of (A.2) is strictly increasing in $m$ regardless of $\delta$. The probability of history $n\mathbf{1}$ is
@@
$\frac{\pi(n\mathbf{1}_{i})}{1-\pi(n\mathbf{1}_{i})}\geq\frac{\pi_{0}}{1-\pi_{0}}\cdot\min_{k\geq n}\frac{(1-\delta)\delta^{k}\cdot(1-\delta(1-\overline{f_{i}}))^{-(k+1)}\cdot\overline{f_{i}}^{k}}{(1-\delta)\delta^{k}\cdot(1-\delta(1-\underline{f_{i}}))^{-(k+1)}\cdot\underline{f_{i}}^{k}}$
$\geq\frac{\pi_{0}}{1-\pi_{0}}\cdot\frac{(1-\delta)\delta^{n}\cdot(1-\delta(1-\overline{f_{i}}))^{-(n+1)}\cdot\overline{f_{i}}^{n}}{(1-\delta)\delta^{n}\cdot(1-\delta(1-\underline{f_{i}}))^{-(n+1)}\cdot\underline{f_{i}}^{n}}=\frac{\pi_{0}}{1-\pi_{0}}\cdot\Big(\frac{1-\delta(1-\underline{f}_{i})}{1-\delta(1-\overline{f}_{i})}\Big)^{n+1}\cdot L_{i}^{n}.$ (A.4)
@@
Appendix B Proof of Lemma 1
  1. 1.
  2. 2.
  3. 3.
Appendix C Proof of Lemma 2
@@
    The monotonicity of the belief map $\pi$ then implies the existence of $\mathbf{h}\leq\mathbf{h^{**}}$ that satisfies $\pi(\mathbf{h})\geq\pi(\mathbf{h^{*}})$. Since history $\mathbf{h^{**}}$ occurs with positive probability in equilibrium, we have $\pi(\mathbf{h^{**}})\geq\pi(\mathbf{h})$ for every $\mathbf{h}\leq\mathbf{h^{**}}$. Therefore, $\pi(\mathbf{h^{**}})\geq\pi(\mathbf{h^{*}})$. However $\mathbf{h}(t)=\mathbf{h^{*}}$ and $\mathbf{h}(\tau)=\mathbf{h^{**}}$ with $t>\tau$ along a realized path of histories, and a history in between them $\mathbf{h}(t-1)=\mathbf{\hat{h}}\neq\mathbf{h^{**}}$. The conclusion that $\pi(\mathbf{h^{**}})\geq\pi(\mathbf{h^{*}})$ contradicts the improvement principle we established earlier, which requires that $\pi(\mathbf{h^{**}})<\pi(\mathbf{h^{*}})$.

### v2

@@ 6 Conclusion
This paper studies the intergenerational transmission of verifiable information in an overlapping generations model that incorporates both physical communication frictions as well as agents’ strategic incentives to conceal unfavorable information. We show that lowering the communication frictions will make the agents more selective in disclosing information. This is because disclosing an additional good signal has two effects on the next agent’s belief: a direct effect which shows the existence of another good signal as well as an indirect effect which shows the existence of another subsequence of bad signals. As communication frictions vanish, the expected length of each bad subsequence increases conditional on each state. Hence, the indirect effect will become stronger, so that fewer signals can have high enough likelihood ratios to overcome the more pronounced indirect effect. The indirect effect is absent in Dye (1985)’s model since any disclosure proves to the receiver that no information is hidden. As a result, an increase in the probability that the sender has information makes his disclosure less selective.
We conclude by discussing several limitations of our analysis. First, as in the classic overlapping generations model of Samuelson (1958), followed by Bhaskar (1998), each agent’s payoff depends only on his action and his immediate successor’s action, [^13] Allowing agents’ payoffs to depend also on their predecessors’ actions, as in Acemoglu and Jackson (2015), will not change our results since even with such a dependence, each agent’s objective is still to maximize their immediate successor’s belief about the state when they choose the set of signals to disclose. but not on future agents’ actions. Suppose each agent $t$’s payoff depends not only on agent $t+1$’s action but also on the action of agent $t+2$’s. When agent $t$ decides which set of signals to disclose, maximizing agent $t+1$’s belief is no longer sufficient since the signals he disclosed also affects agent $t+1$’s ability to influence agent $t+2$’s belief, which in turn affects agent $t$’s payoff. Moreover, aside from the set of signals available to agent $t$, agent $t$’s posterior belief about $\theta$ also affects his incentives to disclose signals since it affects his predictions regarding the realized private signal of agent $t+1$, which together with the signals disclosed by agent $t$, affects agent $t+1$’s disclosure behavior. Solving this dynamic optimization problem is challenging, which we left for future research.
Second, we assume that agent $t$’s payoff takes the form of $u(\theta,a^{t})+v(\theta,a^{t+1})$, that is, agent $t$’s marginal utility from his own action is independent of agent $t+1$’s action. If agent $t$’s payoff takes the general form $w(\theta,a^{t},a^{t+1})$, then as in the previous case, agent $t$ also needs to consider agent $t+1$’s ability to influence agent $t+2$’s belief since agent $t+1$’s incentive to take actions depends not only on his belief about $\theta$ but also on his belief about agent $t+2$’s action. The same challenge arises when solving this dynamic optimization problem.
Third, as in Clark, Fudenberg, and Wolitzky (2021), our main result, Theorem 1, restricts attention to strict steady state equilibria. [^14] Our Theorem 2, which characterizes all constant-threshold equilibria, do not require the strictness of players’ incentives, as indicated in the statement of that result. We do not know whether Theorem 1 holds when we weaken our solution concept from strict monotone equilibria and strict threshold equilibria to pure-strategy monotone equilibria and pure-strategy threshold equilibria. One challenge is that no-turning-back property breaks down when we do not require players to have strict incentives, in which case Lemma 2 as well as the calculations after Lemma 2 and Lemma 3 no longer apply. For example, in threshold equilibria, once the history reaches $\mathbf{h^{*}}$ defined in Section 4, agents may have an incentive to conceal signal $s_{j}$ upon receiving signal $s_{i}$ with $i>j$ if $\pi(\mathbf{h^{*}}[j])=\pi(\mathbf{h^{*}}[k])$ for some $k<j$. In equilibrium, this may in turn boost the next agent’s belief upon observing history $\mathbf{h^{*}}$ as leaving history $\mathbf{h^{*}}$ following bad signals means less adverse selection.
Nevertheless, in the extension we considered where each agent can disclose up to $N$ signals, under generic parameter values $(u,v,\mathbf{f},\pi_{0},\delta)$, all pure-strategy equilibria are strict. This is because when $N$ is finite, there is only a finite number of signal profiles that can be disclosed. If two signal profiles lead to the same expected payoff for the sender, then it leads to a polynomial of $\delta$. Since there is a finite number of signal profiles, there can be at most a finite number of $\delta$ for every $(u,v,\mathbf{f},\pi_{0})$ under which the sender is indifferent.
Appendix A Omitted Proofs
A.1 Proof of Lemma 1
Fix any $\theta$ and pure strategy $\sigma$, [^15] The proof applies to mixed strategies by defining $P^{\sigma}_{\theta}(\mathbf{h},\mathbf{h}^{\prime})=(1-\delta)\mathds{1}_{\mathbf{h}^{\prime}=\mathbf{0}}+\delta\sum_{i=1}^{l}f_{\theta,i}\sigma_{D}(\mathbf{h},s_{i})(\mathbf{h^{\prime}})$. and recall that $\sigma_{D}:H\times S\rightarrow H$ stands for the component of $\sigma$ that maps an agent’s history and private signal to the set of signals disclosed to the next agent. For any $\mathbf{h},\mathbf{h}^{\prime}\in H$, define the probability of moving from history $\mathbf{h}$ to history $\mathbf{h^{\prime}}$ as
$P^{\sigma}_{\theta}(\mathbf{h},\mathbf{h}^{\prime})=(1-\delta)\mathds{1}_{\mathbf{h}^{\prime}=\mathbf{0}}+\delta\sum_{i=1}^{l}f_{\theta,i}\mathds{1}_{\sigma_{D}(\mathbf{h},s_{i})=\mathbf{h^{\prime}}}.$
The above equation defines a stochastic process on histories with transition probability function $P^{\sigma}_{\theta}$. The process is Markov, since each player’s strategy can only depend on their disclosed history and signal, which is i.i.d. over time conditional on the state. Conditional on the selected history being successfully observed by the next generation, we can also define a transition probability $Q^{\sigma}_{\theta}(\mathbf{h},\mathbf{h^{\prime}})=\sum_{i=1}^{l}f_{\theta,i}\mathds{1}_{\sigma_{D}(\mathbf{h},s_{i})=\mathbf{h^{\prime}}}$ and observe that it also defines a Markov process. Thus, we can write
$P^{\sigma}_{\theta}(\mathbf{h},\mathbf{h}^{\prime})=(1-\delta)\mathds{1}_{\mathbf{h}^{\prime}=\mathbf{0}}+\delta Q^{\sigma}_{\theta}(\mathbf{h},\mathbf{h}^{\prime}).$ (A.1)
We show existence and uniqueness of a steady state distribution using the Banach fixed point theorem. Let $T^{\sigma}_{\theta}:\Delta(H)\rightarrow\Delta(H)$ be the transition operator implied by $P^{\sigma}_{\theta}$, so that $(T^{\sigma}_{\theta}(\mu))(\mathbf{h^{\prime}})=\sum_{\mathbf{h}\in H}\mu(\mathbf{h})P^{\sigma}_{\theta}(\mathbf{h},\mathbf{h^{\prime}})$. We claim that this is a contraction mapping when equipped with the total variation distance.
Note that applying only the $Q^{\sigma}_{\theta}$ transition does not increase total variation distance. This follows because $Q^{\sigma}_{\theta}$ is a Markov transition probability. Formally, for any $\mu,\mu^{\prime}\in\Delta(H)$,
$\Big\|\sum_{\mathbf{h}}\mu(\mathbf{h})Q^{\sigma}_{\theta}(\mathbf{h},\cdot)-\sum_{\mathbf{h}}\mu^{\prime}(\mathbf{h})Q^{\sigma}_{\theta}(\mathbf{h},\cdot)\Big\|_{\mathrm{TV}}$ $=\frac{1}{2}\sum_{\mathbf{h}^{\prime}\in H}\Big|\sum_{\mathbf{h}\in H}\big(\mu(\mathbf{h})-\mu^{\prime}(\mathbf{h})\big)Q^{\sigma}_{\theta}(\mathbf{h},\mathbf{h}^{\prime})\Big|$
$\leq\frac{1}{2}\sum_{\mathbf{h}^{\prime}\in H}\sum_{\mathbf{h}\in H}\big|\mu(\mathbf{h})-\mu^{\prime}(\mathbf{h})\big|\,Q^{\sigma}_{\theta}(\mathbf{h},\mathbf{h}^{\prime})$
$=\frac{1}{2}\sum_{\mathbf{h}\in H}\big|\mu(\mathbf{h})-\mu^{\prime}(\mathbf{h})\big|\sum_{\mathbf{h}^{\prime}\in H}Q^{\sigma}_{\theta}(\mathbf{h},\mathbf{h}^{\prime})$
$=\frac{1}{2}\sum_{\mathbf{h}\in H}\big|\mu(\mathbf{h})-\mu^{\prime}(\mathbf{h})\big|$
$=\|\mu-\mu^{\prime}\|_{\mathrm{TV}},$
where we used that $\sum_{\mathbf{h}^{\prime}\in H}Q^{\sigma}_{\theta}(\mathbf{h},\mathbf{h}^{\prime})=1$ for each $\mathbf{h}$.
By (A.1) and the above inequality, we have that for any $\mu,\mu^{\prime}\in\Delta(H)$,
$\|T^{\sigma}_{\theta}(\mu)-T^{\sigma}_{\theta}(\mu^{\prime})\|_{\mathrm{TV}}\leq\delta\,\|\mu-\mu^{\prime}\|_{\mathrm{TV}}.$
Since $\delta\in(0,1)$, $T^{\sigma}_{\theta}$ is a contraction on $(\Delta(H),\|\cdot\|_{\mathrm{TV}})$. By the Banach fixed point theorem, there exists a unique steady state distribution over histories $\mu_{\theta}\in\Delta(H)$.
A.2 Proof of Proposition 1
@@
If $L_{i}>1$, then the value of (A.3) is strictly increasing in $m$ regardless of $\delta$. The probability of history $n\mathbf{1}_{i}$ is
@@
$\frac{\pi(n\mathbf{1}_{i})}{1-\pi(n\mathbf{1}_{i})}$ $\geq\frac{\pi_{0}}{1-\pi_{0}}\cdot\min_{k\geq n}\Big\{\frac{(1-\delta)\,\delta^{k}\big(1-\delta(1-\overline{f}_{i})\big)^{-(k+1)}\overline{f}_{i}^{\,k}}{(1-\delta)\,\delta^{k}\big(1-\delta(1-\underline{f}_{i})\big)^{-(k+1)}\underline{f}_{i}^{\,k}}\Big\}$
$\geq\frac{\pi_{0}}{1-\pi_{0}}\cdot\frac{(1-\delta)\,\delta^{n}\big(1-\delta(1-\overline{f}_{i})\big)^{-(n+1)}\overline{f}_{i}^{\,n}}{(1-\delta)\,\delta^{n}\big(1-\delta(1-\underline{f}_{i})\big)^{-(n+1)}\underline{f}_{i}^{\,n}}$
$=\frac{\pi_{0}}{1-\pi_{0}}\cdot\left(\frac{1-\delta(1-\underline{f}_{i})}{1-\delta(1-\overline{f}_{i})}\right)^{\!n+1}L_{i}^{\,n}.$ (A.5)
@@
A.3 Proof of Proposition 2
For any $n\geq 2$, fix $\delta^{\prime}<\delta$ so that $\delta^{\prime},\delta\in(\delta_{n+1},\delta_{n}]$. We first show that $I(\theta;\mathbf{h}_{\delta^{\prime}})\leq I(\theta;\mathbf{h}_{\delta})$. Observe that both $\mathbf{h}_{\delta}$ and $\mathbf{h}_{\delta^{\prime}}$ involve disclosing signals $\{s_{1},...,s_{n}\}$. We refer to the sequence of signals after the last communication failure (including the ones concealed) as the complete history of game. Since the expected length of the complete history in the steady state equals $\delta/(1-\delta)$, the expected number of signals disclosed is increasing in $\delta$. Hence, if $\delta^{\prime}<\delta$, then the random variable $\mathbf{h}_{\delta^{\prime}}$ can be represented as a garbling of the random variable $\mathbf{h}_{\delta}$, which implies that $I(\theta;\mathbf{h}_{\delta^{\prime}})\leq I(\theta;\mathbf{h}_{\delta})$. Observe also that for such $\delta$, $I(\theta;\mathbf{h}_{\delta})$ varies continuously since $\delta/(1-\delta)$ is continuous in $\delta$.
Fix any cutoff $\delta_{n}$. By definition, signals $\{s_{1},...,s_{n}\}$ are disclosed in the most informative constant-threshold equilibrium when $\delta$ is slightly below $\delta_{n}$ and signals $\{s_{1},...,s_{n-1}\}$ are disclosed in the most informative constant-threshold equilibrium when $\delta$ is slightly above $\delta_{n}$ At the cutoff $\delta_{n}$, the incentive constraint to disclose signal $n$ is binding, so that by equation (3.8) we have
But because of the binding constraint, disclosing an additional realization of $s_{n}$ does not change the likelihood ratio of the public belief. Hence, when $\delta=\delta_{n}$, we have $\pi(h_{1},...,h_{n-1},0,...,0)=\pi(h_{1},...,h_{n-1},h_{n},0,...,0)$ for any $h_{n}\in\mathbb{N}$ and $(h_{1},...,h_{n-1})\in\mathbb{N}^{n-1}$. Therefore, we have $I(\theta;\mathbf{h}^{(n)}_{\delta_{n}})=I(\theta;\mathbf{h}^{(n-1)}_{\delta_{n}})=I(\theta;\mathbf{h}_{\delta_{n}})$ and $I(\theta;\mathbf{h}_{\delta})$ is continuous at $\delta=\delta_{n}$.
A.4 Proof of Lemma 2
@@
    The monotonicity of the belief map $\pi$ then implies the existence of $\mathbf{h}\leq\mathbf{h^{**}}$ such that $\mathbf{h}$ and $\mathbf{h^{*}}$ share the same length and moreover, $\pi(\mathbf{h})\geq\pi(\mathbf{h^{*}})$. Since history $\mathbf{h^{**}}$ occurs with positive probability in equilibrium, we have $\pi(\mathbf{h^{**}})\geq\pi(\mathbf{h})$ for every $\mathbf{h}\leq\mathbf{h^{**}}$. Therefore, $\pi(\mathbf{h^{**}})\geq\pi(\mathbf{h^{*}})$. However $\mathbf{h}(t)=\mathbf{h^{*}}$ and $\mathbf{h}(\tau)=\mathbf{h^{**}}$ with $t>\tau$ along a realized path of histories, and a history in between them $\mathbf{h}(t-1)=\mathbf{\hat{h}}\neq\mathbf{h^{**}}$. The conclusion that $\pi(\mathbf{h^{**}})\geq\pi(\mathbf{h^{*}})$ contradicts the improvement principle we established earlier, which requires that $\pi(\mathbf{h^{**}})<\pi(\mathbf{h^{*}})$.
Remark:
The proof of Lemma 2 extends to the model discussed in Section 5.3 where communication success rate weakly decreases as the disclosed history gets longer. To see this, let us revisit the three cases considered in the proof. The first two cases only involve comparison between beliefs at two histories of the same length, namely, $\mathbf{h^{*}}$ and $\mathbf{h^{*}}-\mathbf{1}_{j}+\mathbf{1}_{i}$. Since the communication success rate depends only on the length of the disclosed history, an agent prefers to disclose $\mathbf{h^{*}}$ to $\mathbf{h^{*}}-\mathbf{1}_{j}+\mathbf{1}_{i}$ if and only if $\pi(\mathbf{h^{*}})\geq\pi(\mathbf{h^{*}}-\mathbf{1}_{j}+\mathbf{1}_{i})$. In the third case, recall the construction of history $\mathbf{h^{**}}$ as well as the existence of history $\mathbf{h}$ such that $\mathbf{h}\leq\mathbf{h^{**}}$, $\mathbf{h}$ and $\mathbf{h^{*}}$ share the same length, and $\pi(\mathbf{h})\geq\pi(\mathbf{h^{*}})$. Therefore, the expected payoff from disclosing $\mathbf{h}$ is weakly greater than the expected payoff from disclosing $\mathbf{h^{*}}$, as the communication success rates are the same yet disclosing $\mathbf{h}$ leads to a weakly higher belief. This leads to a contradiction since disclosing $\mathbf{h}$ is available when the history is $\mathbf{h^{**}}$ yet history $\mathbf{h^{*}}$ is reached after $\mathbf{h^{**}}$ before the next time communication breaks down.
A.5 Proof of Lemma 3
  1. 1.
  2. 2.
  3. 3.
A.6 Proof of Proposition 3
If the sender uses Strategy $k$, then the receiver will only observe $\mathbf{h}\equiv(k+1)\mathbf{1}_{1}+\mathbf{1}_{2}$ and $j\mathbf{1}_{1}$ for every $j\neq k+1$ with positive probability. The receiver’s posterior beliefs, given by the mapping $\pi:H\rightarrow[0,1]$, are pinned down by Bayes rule following these positive probability events. Therefore, in order to verify that Strategy $k$ is part of a strict equilibrium, we only need to verify that the following three conditions are satisfied:
  1. 1.
    $\pi(j\mathbf{1}_{1})$ is strictly increasing in $j$ for every $j\neq k+1$.
  2. 2.
    $\pi((k+2)\mathbf{1}_{1})>\pi(\mathbf{h})$.
  3. 3.
    $\pi(\mathbf{h})>\pi(k\mathbf{1}_{1})$.
The first two conditions are satisfied for all $k\in\mathbb{N}$, as implied by expressions (3.2) to (3.5) in the proof of Theorem 2. In what follows, we show that for every $\delta\in(0,1)$, there exists a large enough $k$ such that the third condition is also satisfied. Applying (3.2) and (3.4), we know that if the sender uses Strategy $k$ in the one-shot game, then conditional on the state being $\theta$, the receiver will observe $k\mathbf{1}_{1}$ with probability
$(1-\delta)\delta^{k}\Big(1-\delta(1-f_{\theta,1})\Big)^{-(k+1)}f_{\theta,1}^{k}+(1-\delta)\delta^{k+1}\Big(1-\delta(1-f_{\theta,1}-f_{\theta,2})\Big)^{-(k+2)}f_{\theta,1}^{k+1},$
where the first term is the probability that the sender’s observes $k$ signal $s_{1}$ and an unknown number of signals from $s_{2}$ to $s_{l}$, and the second term is the probability that the sender observes $k+1$ signal $s_{1}$, no signal $s_{2}$, as well as an unknown number of signals from $s_{3}$ to $s_{l}$. Under the same sender-strategy and conditional on state $\theta$, the receiver will observe $\mathbf{h}$ with probability
$(1-\delta)\delta^{k+1}\Big(1-\delta(1-f_{\theta,1})\Big)^{-(k+2)}f_{\theta,1}^{k+1}-(1-\delta)\delta^{k+1}\Big(1-\delta(1-f_{\theta,1}-f_{\theta,2})\Big)^{-(k+2)}f_{\theta,1}^{k+1},$
where the first term is the probability that the sender observes $k+1$ signal $s_{1}$ and an unknown number of signals from $s_{2}$ to $s_{l}$ and the second term coincides with the second term in the previous expression except with the opposite sign. According to Bayes rule, we know that $\pi(\mathbf{h})>\pi(k\mathbf{1}_{1})$ if and only if the value of
$\frac{\Big(1-\delta(1-f_{\theta,1})\Big)^{-(k+1)}+\delta f_{\theta,1}\Big(1-\delta(1-f_{\theta,1}-f_{\theta,2})\Big)^{-(k+2)}}{f_{\theta,1}\Big(1-\delta(1-f_{\theta,1})\Big)^{-(k+2)}-f_{\theta,1}\Big(1-\delta(1-f_{\theta,1}-f_{\theta,2})\Big)^{-(k+2)}}$ (A.6)
strictly decreases once $\theta$ increases from $\underline{\theta}$ to $\overline{\theta}$. This is the case if and only if
$\frac{\Big(\overline{X}^{-(k+1)}+\delta\overline{f_{1}}\cdot\overline{Y}^{-(k+2)}\Big)\cdot\Big(\underline{X}^{-(k+2)}-\underline{Y}^{-(k+2)}\Big)}{\Big(\underline{X}^{-(k+1)}+\delta\underline{f_{1}}\cdot\underline{Y}^{-(k+2)}\Big)\cdot\Big(\overline{X}^{-(k+2)}-\overline{Y}^{-(k+2)}\Big)}<\frac{\overline{f_{1}}}{\underline{f_{1}}}=L_{1},$ (A.7)
where
$\overline{X}\equiv 1-\delta(1-\overline{f_{1}}),\quad\underline{X}\equiv 1-\delta(1-\underline{f_{1}})$
[…]
Let
[…]
Inequality (A.7) can then be rewritten as
$\frac{(\overline{X}+\delta\overline{f_{1}}\cdot\overline{R}^{k+2})(1-\underline{R}^{k+2})}{(\underline{X}+\delta\underline{f_{1}}\cdot\underline{R}^{k+2})(1-\overline{R}^{k+2})}<L_{1}$ (A.8)
For every $\delta\in(0,1)$, we know that $\overline{X}<\overline{Y}$ and $\underline{X}<\underline{Y}$, which implies that $0<\underline{R}<\overline{R}<1$. Therefore, fix any $\delta\in(0,1)$, both $\overline{R}^{k+2}$ and $\underline{R}^{k+2}$ converge to 0 as $k\rightarrow+\infty$. This suggests that left-hand-side of (A.8) converges to $\overline{X}/\underline{X}$ as $k\rightarrow+\infty$. Since $L_{1}>\overline{X}/\underline{X}$ for any $\delta\in(0,1)$, we know that for any $\delta\in(0,1)$, there exists a large enough $k$ under which inequality (A.8) is satisfied. This verifies that no matter how close $\delta$ is to $1$, there exists a strict monotone equilibrium under which signal $s_{2}$ is disclosed with positive probability.
A.7 Proof of Proposition 4
Suppose by way of contradiction that for every $\delta$ close enough to $1$, there exists a strict equilibrium in the one-shot disclosure game where the sender uses a threshold strategy $\sigma_{D}$ under which there exists $j\geq 2$ such that the $j$th entry of $\sigma_{D}(\mathbf{h})$ is strictly positive for some $\mathbf{h}\in H$.
Fix any such equilibrium and without loss of generality, let $j\in\{2,...,l\}$ denote the largest integer such that there exists $\mathbf{h}$ such that the $j$th entry of $\sigma_{D}(\mathbf{h})$ is not zero. Let $\mathbf{h^{*}}\equiv(h_{1}^{*},...,h_{l}^{*})\in H$ denote a history at which $s_{j}$ is disclosed. Since $\sigma_{D}$ is a threshold strategy, we have $\sigma_{D}(\mathbf{h^{*}})=\mathbf{h^{*}}[j]$. For every $k\in\mathbb{N}$, let $\mathbf{h^{k}}\equiv\mathbf{h^{*}}[j]+k\mathbf{1}_{j}$.
  Lemma 4.
For every $\mathbf{h}\in H$ such that there exists an integer $k\geq 1$ that satisfies $\mathbf{h}[j]=\mathbf{h^{k}}$, we have $\sigma_{D}(\mathbf{h})=\mathbf{h^{k}}[j]$.
Proof of Lemma 4:.
Since $j$ is the highest signal index that can be disclosed, the equilibrium is strict, and $\sigma_{D}$ is a threshold strategy, we know that $\sigma_{D}(\mathbf{h^{*}}[j])=\mathbf{h^{*}}[j]$ and
$\pi(\mathbf{h^{*}}[j])>\pi(\mathbf{h})\textrm{ for every }\mathbf{h}<\mathbf{h^{*}}[j].$ (A.9)
Once the sender observes $\mathbf{h^{1}}$, we from (A.9) that $\sigma_{D}(\mathbf{h^{1}})$ cannot be anything strictly less than $\mathbf{h^{*}}[j]$ since disclosing $\mathbf{h^{*}}[j]$ is available. Since $\sigma_{D}$ is a threshold strategy, all available signal $s_{j}$ must be disclosed as long as one $s_{j}$ is disclosed, which implies that $\sigma_{D}(\mathbf{h^{1}})=\mathbf{h^{1}}[j]$. Using the same logic, one can show that for every $k\in\mathbb{N}$, the hypothesis that $\sigma_{D}(\mathbf{h^{k}})=\mathbf{h^{k}}[j]$ implies that $\sigma_{D}(\mathbf{h^{k+1}})=\mathbf{h^{k+1}}[j]$. This induction argument implies that $\sigma_{D}(\mathbf{h^{k}})=\mathbf{h^{k}}[j]$ for every $k\geq 1$. The definition of $j$ implies that signals that do not belong to $\{s_{1},...,s_{j}\}$ will never be disclosed under $\sigma_{D}$, which implies that $\sigma_{D}(\mathbf{h})=\mathbf{h^{k}}[j]$ for every $\mathbf{h}$ that satisfies $\mathbf{h}[j]=\mathbf{h^{k}}$. ∎
Lemma 4 suggests that if the receiver observes history $\mathbf{h^{k}}$, then she infers that the sender has disclosed all the available $\{s_{1},...,s_{j}\}$ and has concealed all the other signals. Recall our hypothesis that $j\geq 2$. Applying the calculations in the proof of Theorem 2 to $\pi(\mathbf{h^{1}})$ and $\pi(\mathbf{h^{2}})$, we know that there exists $\delta^{*}\in(0,1)$ such that $\pi(\mathbf{h^{1}})>\pi(\mathbf{h^{2}})$ as long as $\delta>\delta^{*}$. This contradicts the sender’s incentive to disclose $\mathbf{h^{2}}$ at $\mathbf{h^{2}}$ which requires that $\pi(\mathbf{h^{1}})\leq\pi(\mathbf{h^{2}})$.